# Local spin polarization by color-field correlators and momentum anisotropy

### v2

@@ Abstract
Spin polarization from color-field correlators.— Before introducing QKT, we provide an intuitive argument for generating spin polarization from color fields. We consider the chromo-Lorentz force, ${\bm{F}}^{a}=g({\bm{E}}^{a}+{\bm{u}}\times{{\bm{B}}^{a}})$, and the color-octet polarization pseudo-vector from chromo-magnetic polarization and chromo-spin Hall effect, ${\bm{P}}^{a}=g({\bm{B}}^{a}-{\bm{u}}\times{{\bm{E}}^{a}})$, for the quark transport, where $\bm{u}$} as a flow velocity delineates momentum anisotropy. The color-singlet polarization pseudo-vector for quarks with momentum $\bm{p}$} can accordingly be constructed by the correlation, $\langle({\bm{p}}\cdot{\bm{F}}^{a}){\bm{P}}^{a}\rangle\sim({\bm{p}}\times{\bm{u}})(\langle{\bm{B}}^{a}\cdot{\bm{B}}^{a}\rangle+\langle{\bm{E}}^{a}\cdot{\bm{E}}^{a}\rangle)$, when having just non-vanishing parity-even color-field correlators. We now derive this effect in the QKT.
@@
$\mathcal{D}_{p\nu}\equiv\partial_{p\nu}p_{0}^{-1}-\epsilon_{\bm{p}}\partial_{p_{\perp}\nu}\epsilon_{\bm{p}}^{-2}+p_{\perp\nu}\epsilon_{\bm{p}}^{-2}\partial_{p0}$ (7)
and we have introduced $\langle\,\rangle$ to denote the ensemble average over color sources. For a more detailed derivation we refer to Refs. [68, 69].
@@
eventually approaching thermal equilibrium. We then evaluate the LSP via
@@
We have numerically evaluated Eq. (19) by integrating over $r_{m}-\Delta t\leq r\leq r_{m}$, $0\leq\Phi\leq 2\pi$, and $1\leq\eta\leq 1$ at a fixed $\tilde{\tau}_{f}$, where the regime $r_{m}-\Delta t\leq r\leq r_{m}$ is set to capture only the corona contribution. Here we focus on the case with $\tilde{\tau}_{f}\approx 0$ such that $\tau_{f}\approx r\leq r_{m}$. We take $\hbar=1$, $N_{c}=3$, $m=0.5$ GeV as the constituent strange quark mass, $M_{\Lambda}=1.12$ GeV, $\Delta t=0.2$ fm and $r_{m}\approx\tau^{\rm max}_{f}=0.5$ fm as the typical time scale smaller than $1$ fm for glasma phase. We set $\delta=0.3$ by analogy with the anisotropy parameter adopted in the thermal model for non-central collisions [90, 79], while we take $u_{T}=0.01$ and $u_{T}=0.2$ for Eq. (17) and Eq. (18), respectively, as the weak non-equilibrium flow gradually generated by pressure gradients of glasma. The $\phi$ dependence of $\mathcal{P}^{z}$ at fixed $p_{T}=2$ GeV for $Q_{s}=1.5$ or $2$ GeV are shown in Fig. 2 and Fig. 3, where the $\sin 2\phi$ structure is observed as anticipated. In addition, $\mathcal{P}^{z}$ with transverse-momentum dependence at fixed $\phi=\pi/4$ is also presented therein. For large $p_{T}$, the $G_{1}$ contribution becomes more substantial when $|{\bm{p}}\cdot{\bm{u}}|\sim Q_{s}$, which may cause the sign flipping of $\mathcal{P}^{z}$ at sufficiently large $p_{T}$.
@@
where $N$ denotes the particle number and the subscripts $\rm GL$ and $\rm QGP$ represent the contributions from the glasma (corona) and QGP (core), respectively. In general, $N_{\rm GL}$ and $N_{\rm QGP}$ should be more dominant at high $p_{T}$ for small systems and at low $p_{T}$ for large systems, respectively. More sophisticated modeling and simulations for the core and corona scenario similar to Ref. [92] will be needed to quantify the results.
Recent measured strength of $\mathcal{P}^{z}$ in proton-nucleus collisions at LHC shows a mild opposite trend against multiplicity [22], which could be qualitatively explained by the dominant contribution from the corona in smaller collision systems due to the suppression of $N_{\rm QGP}$ with lower multiplicity. When the local equilibrium condition for spin degrees of freedom is reached, there exist additional contributions from thermal vorticity and thermal-shear corrections for $\mathcal{P}^{z}_{\rm QGP}$, further competing with the glasma effect. Finally, we remark that the initial flow for quarks in glasma could also stem from mini-jets and thus the momentum anisotropy and hydrodynamic flow may be misaligned with increasing multiplicity.

### v3

@@ Abstract
Spin polarization from color-field correlators.— Before introducing QKT, we provide an intuitive argument for generating spin polarization from color fields. We consider the chromo-Lorentz force, ${\bm{F}}^{a}=g({\bm{E}}^{a}+{\bm{u}}\times{{\bm{B}}^{a}})$, and the color-octet polarization pseudo-vector from chromo-magnetic polarization and chromo-spin Hall effect, ${\bm{P}}^{a}=g({\bm{B}}^{a}-{\bm{u}}\times{{\bm{E}}^{a}})$, for the quark transport, where $\bm{u}$} as a flow velocity delineates momentum anisotropy. The color-singlet polarization pseudo-vector for quarks with momentum $\bm{p}$} can accordingly be constructed by the correlation, $\langle({\bm{p}}\cdot{\bm{F}}^{a}){\bm{P}}^{a}\rangle\sim({\bm{p}}\times{\bm{u}})(\langle{\bm{B}}^{a}\cdot{\bm{B}}^{a}\rangle+\langle{\bm{E}}^{a}\cdot{\bm{E}}^{a}\rangle)$, when having just non-vanishing parity-even color-field correlators. Here we have introduced $\langle\,\rangle$ to denote the ensemble average over color sources. We now derive this effect in the QKT.
@@
$\mathcal{D}_{p\nu}\equiv\partial_{p\nu}p_{0}^{-1}-\epsilon_{\bm{p}}\partial_{p_{\perp}\nu}\epsilon_{\bm{p}}^{-2}+p_{\perp\nu}\epsilon_{\bm{p}}^{-2}\partial_{p0}.$ (7)
For a more detailed derivation we refer to Refs. [68, 69].
@@
eventually approaching thermal equilibrium characterized by $Q_{s}\rightarrow T$ and $c\rightarrow 1$. Here we set $0\leq c\leq 1$ to describe the overpopulation of quarks driven by the high density of gluons at low energy while satisfying the Pauli exclusion principle. We then evaluate the LSP via
@@
We have numerically evaluated Eq. (19) by integrating over $r_{m}-\Delta t\leq r\leq r_{m}$, $0\leq\Phi\leq 2\pi$, and $1\leq\eta\leq 1$ at a fixed $\tilde{\tau}_{f}$, where the regime $r_{m}-\Delta t\leq r\leq r_{m}$ is set to capture only the corona contribution. Here we focus on the case with $\tilde{\tau}_{f}\approx 0$ such that $\tau_{f}\approx r\leq r_{m}$. We take $\hbar=1$, $N_{c}=3$, $m=0.5$ GeV as the constituent strange quark mass, $M_{\Lambda}=1.12$ GeV, $\Delta t=0.2$ fm and $r_{m}\approx\tau^{\rm max}_{f}=0.5$ fm as the typical time scale smaller than $1$ fm for glasma phase. We set $\delta=0.3$ by analogy with the anisotropy parameter adopted in the thermal model for non-central collisions [90, 79], while we take $u_{T}=0.01$ and $u_{T}=0.2$ for Eq. (17) and Eq. (18) characterizing the initial and final quark distributions in the glasma phase, respectively, as the weak non-equilibrium flow gradually generated by pressure gradients of glasma in time. The $\phi$ dependence of $\mathcal{P}^{z}$ at fixed $p_{T}=2$ GeV for $Q_{s}=1.5$ or $2$ GeV are shown in Fig. 2 and Fig. 3, where the $\sin 2\phi$ structure is observed as anticipated. In addition, $\mathcal{P}^{z}$ with transverse-momentum dependence at fixed $\phi=\pi/4$ is also presented therein. For large $p_{T}$, the $G_{1}$ contribution becomes more substantial when $|{\bm{p}}\cdot{\bm{u}}|\sim Q_{s}$, which may cause the sign flipping of $\mathcal{P}^{z}$ at sufficiently large $p_{T}$.
@@
where $N$ denotes the particle number and the subscripts $\rm GL$ and $\rm QGP$ represent the contributions from the glasma (corona) and QGP (core), respectively. In general, $N_{\rm GL}$ and $N_{\rm QGP}$ should be more dominant at high $p_{T}$ for small systems and at low $p_{T}$ for large systems, respectively. More sophisticated modeling and simulations for the core and corona scenario similar to Ref. [92] will be needed to quantify the results. Despite the dependence on transverse momenta and centrality, it is found therein that the corona contribution for hadron yields is about 50% when the multiplicity per rapidity reaches $\mathcal{O}(10^{1})$ for both p+p and Pb+Pb collisions. For a heuristic estimation, by taking $\mathcal{P}^{z}_{\rm GL}=4\times 10^{-3}$ and $\mathcal{P}^{z}_{\rm QGP}=-1.5\times 10^{-3}$ from Fig. 2 and Fig. 4, Eq. (22) still yields $\mathcal{P}^{z}\approx 0.1\%$ for multiplicity around tens to a few hundreds as for small collision systems.
Recent measured strength of $\mathcal{P}^{z}$ in proton-nucleus collisions at LHC shows a mild opposite trend against multiplicity [22], which could be qualitatively explained by the substantial contribution from the corona in smaller collision systems due to the suppression of $N_{\rm QGP}$ with lower multiplicity. When the local equilibrium condition for spin degrees of freedom is reached, there exist additional contributions from thermal vorticity and thermal-shear corrections for $\mathcal{P}^{z}_{\rm QGP}$, further competing with the glasma effect. Finally, we remark that the initial flow for quarks in glasma could also stem from mini-jets and thus the momentum anisotropy and hydrodynamic flow may be misaligned with increasing multiplicity.